An importer of Swiss watches has an account payable of CHF750,000 due in 90 days. The following data is available:

Rates and prices in US-cents/CHF:

Spot rate: 71.42 cents/CHF

90-day forward rate: 71.14 cents/CHF

US –dollar 90-day interest rate: 3.75% per year

Swiss franc 90-day interest rate: 5.33% per year

Option Data in cents/CHF

Strike Call Put

70 2.55 1.42

72 1.55 2.40

a) Assess the USD cost to the importer in 90 days if it uses a call option to hedge its CHF750,000 account payable. Use the call with a strike price of 72 cents/CHF and include the option call premium in the cost.

b) What will be the cost of the payable in 90 days if a forward contract is used?

c) By how much must the CHF weaken relative to the USD, from 71.42 cents/CHF before the call option provides a lower cost than the forward hedge?
Answer #1

Part (a)

The USD cost to the importer in 90 days if it uses a call option to hedge its CHF750,000 account payable = (K + C) x Amount payable = (72 + 1.55) cents / CHF x CHF750,000 = $ 551,625.00

Part (b)

the cost of the payable in 90 days if a forward contract is used = Forward rate x Amount payable = 71.14 cents/CHF x CHF 750,000 = $ 533,550.00

Part (c)

S + C = F

Hence, S = F - C = 71.14 - 1.55 = Cents 69.59 / CHF

Hence, implied weakening of CHF = S0 - S = 71.42 - 69.59 = Cents 1.83 / CHF

Hence, the CHF must weaken by Cents 1.83 / CHF, relative to the USD, from 71.42 cents/CHF before the call option provides a lower cost than the forward hedge